OpenStudy (anonymous):

Geometry Help! Please if m(angle)AXC=8x - 7 and m(angle)AXB = 3x + 10, find m(angle)AXC

OpenStudy (anonymous):



OpenStudy (anonymous):

Angle AXC is bisected by XB

OpenStudy (bibby):

By definition, AXB is half of AXC

OpenStudy (anonymous):

Yess..

OpenStudy (bibby):

\(\huge \frac{8x - 7 }{2}= 3x + 10\)

OpenStudy (anonymous):

Soo.. the answer isn't going to be a single number?

OpenStudy (bibby):

I did the math, AXC is a whole number as far as I can tell

OpenStudy (anonymous):

oh, you have to solve it, sorry

OpenStudy (bibby):

I don't have to do anything. I set the equation up. Solve for x.

OpenStudy (anonymous):

I know, I was referring like, "in order to solve the problem, you (like I would need to) need to solve for x"

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

 It also says ray XA and XE are opposite rays. if m(angle)CXD=4x +6, m(angle)DXE=3x+1, and m(angle)=8x-2, find m(angle)DXE?